# DiffOG: Differentiable Policy Trajectory Optimization with Generalizability

### v4

@@ IV-C Ablation Study on Static $\mathbf{Q}$, Matrix Learning, and Transformer
From equation (9), we can observe that if we set $\mathbf{Q}=-\mathbf{I}$, the term $\bigl{\|}\mathbf{R}\mathbf{y}_{t}-\mathbf{g}_{t}\bigr{\|}^{2}$ in the optimization problem (8) simplifies to $\bigl{\|}\mathbf{y}_{t}-\mathbf{a}_{t}\bigr{\|}^{2}$, because $\mathbf{Q}=\mathbf{R}^{\mathrm{T}}\mathbf{R}$ and $\mathbf{g}_{t}=-(\mathbf{R}^{\mathrm{T}})^{-1}\mathbf{a}_{t}$. Under this setting, (8) becomes a trajectory optimization problem with both a smoothing cost and a distance cost. On the other hand, instead of using a transformer as in DiffOG, we can also directly learn the matrix $\mathbf{Q}$. In this section, we conduct experiments on the following three different cases:
@@
Additionally, during refine training, due to the presence of a multi-modal action distribution in demonstration dataset [5, 7], the actions generated by the pre-trained policy may differ from the targets in the dataset. This conflict between the targets and inputs does not exist in dataset training paradigm of DiffOG and matrix learning. From the results of matrix-learning refine and DiffOG refine, we can observe that trajectory optimization based on matrix learning struggles to handle such conflict. However, DiffOG can handle this conflict with transformer encoder.
@@
TABLE VI: We report the batched average inference time across three simulated tasks: lift, can, and square. Each measurement is based on a batch size of 25 and averaged over 50 inference runs. All methods are evaluated using 100 DDPM steps [9].
@@
[…]
@@ IV-D Comparison with a Constrained Visuomotor Policy
We conducted comparison experiments between DiffOG and constrained visuomotor policy Leto [33] on 7 tasks. The results are shown in Table VII. The results in the table compare DiffOG and Leto under the same constraint bounds.
@@
TABLE VII: Success rate of benchmarking DiffOG with Leto [33].
@@ V Limitations and Future Work
Moreover, our method has demonstrated strong empirical performance under widely adopted action representations in imitation learning, such as joint angles [36, 29] and 6D end-effector velocities [13, 27, 37]. Beyond these settings, we identify several promising directions for future work: 1) Certain action spaces that are less common in current imitation learning benchmarks, such as joint torques, also conform to the formulations of DiffOG. While not yet widely used, joint torques hold strong potential for force-based imitation learning, as they provide direct control over physical interactions. Extending our method to torque-based action spaces could enable more contact-rich manipulation. 2) Future extensions of the framework may address more complex representations, such as absolute $SO(3)$ rotations. In such cases, the assumption of simple discrete-time derivatives may no longer hold, requiring more expressive formulations of the underlying optimization problem.
@@ VI Conclusion
We introduced DiffOG, a learning-based trajectory optimization framework that combines the strengths of neural networks with an interpretable differentiable optimization layer. By integrating trajectory smoothness, adherence to hard constraints, and fidelity to demonstrations into the training objectives, DiffOG enables visuomotor policies to produce smoother and safer action trajectories while maintaining performance. Through extensive evaluations on 11 simulated tasks and 2 real-world tasks, we demonstrated that DiffOG consistently outperforms existing trajectory processing baselines, such as greedy constraint clipping and penalty-based trajectory optimization, as well as constrained visuomotor policy Leto [33]. Our results demonstrate the generalization capability of DiffOG across diverse tasks.
@@ VI-B Action Space Assumption
1. The action space consists of joint angles, where the derivative corresponds to joint velocities. This type of action space is commonly used in systems that collect data through leader-follower arms, such as ALOHA [36] and GELLO [29].
2. The action space consists of the robot end-effector linear position in $xyz$, where the derivative corresponds to linear velocity. This action space is used in the push-T task in [5] and the Meta-World benchmark [34].
Another widely adopted action space is $[v_{x},\,v_{y},\,v_{z},\,\dot{r}_{x},\,\dot{r}_{y},\,\dot{r}_{z}]^{\mathrm{T}}\in\mathbb{R}^{6},$ where $\mathbf{v}_{xyz}=[v_{x},v_{y},v_{z}]^{\mathrm{T}}$ is translational velocity, and $\dot{\mathbf{r}}=[\dot{r}_{x},\dot{r}_{y},\dot{r}_{z}]^{\mathrm{T}}$ is the derivative of the axis-angle rotation vector [13, 27, 37, 33]. Its time derivative, $[\dot{v}_{x},\,\dot{v}_{y},\,\dot{v}_{z},\,\ddot{r}_{x},\,\ddot{r}_{y},\,\ddot{r}_{z}]^{\!\mathrm{T}},$ includes linear accelerations and axis-angle second derivatives, which differ from the true angular acceleration $\dot{\bm{\omega}}$.
Although $\ddot{\mathbf{r}}$ does not directly represent $\dot{\bm{\omega}}$, constraining $\ddot{\mathbf{r}}$ effectively regulates $\dot{\bm{\omega}}$. The angular acceleration is related by $\dot{\bm{\omega}}=\mathbf{J}(\mathbf{r})\ddot{\mathbf{r}}+\dot{\mathbf{J}}(\mathbf{r})\dot{\mathbf{r}}$, where $\mathbf{J}(\mathbf{r})$ is the Jacobian matrix mapping axis-angle velocity to angular velocity. In practice, $\mathbf{J}(\mathbf{r})$ typically remains bounded, as evidenced by empirical successes of such action space in imitation learning frameworks [13, 27, 37], which achieve stable rotations even without explicit constraints. Different from these prior work, DiffOG explicitly and effectively constrains rotational motion, resulting in smoother and more controlled behaviors. Extensive experimental results demonstrate the effectiveness of DiffOG.
@@ VI-C Task Descriptions and Experimental Setup
For the end-effector 6-DOF velocity action space, we include five tasks: lift, can, square, tool hang, and transport, all sourced from Robomimic benchmark [13]. For the end-effector linear position action space, we include six tasks: push-T from [5], along with pick place wall, shelf place, disassemble, stick push, and stick pull from Meta-World [34]. Notably, we selected all the “very hard” tasks identified in Meta-World as marked in [35]. Lastly, for the joint angle action space, we include two real-world tasks, arrange desk and move the stack, with data collection and policy experiments conducted using ALOHA [36].
@@
Lift 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 2×3×84×842$\times$ 3$\times$ 84$\times$ 842 × 3 × 84 × 84
Can 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 2×3×84×842$\times$ 3$\times$ 84$\times$ 842 × 3 × 84 × 84
Square 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 2×3×84×842$\times$ 3$\times$ 84$\times$ 842 × 3 × 84 × 84
Tool Hang 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 2×3×240×2402$\times$ 3$\times$ 240$\times$ 2402 × 3 × 240 × 240
Transport 0.1 0.1 200 12-DOF Vel. + 2-DOF Grasping 0.05 s 4 4×3×84×844$\times$ 3$\times$ 84$\times$ 844 × 3 × 84 × 84
Push-T 0.05 [12.5,12.2] Pixels 90 2-DOF Position 0.1 s 1 1×3×96×961$\times$ 3$\times$ 96$\times$ 961 × 3 × 96 × 96
Pick Place Wall 0.6 [0.53, 0.92, 3.75] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Shelf Place 0.6 [2.32, 1.47, 4.13] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Disassemble 0.6 [0.65, 1.10, 0.61] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Stick Push 0.6 [1.03, 0.15, 0.68] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Stick Pull 0.6 [4.46, 1.44, 2.49] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 3×5123$\times$ 5123 × 512
Arrange Desk 0.1 [0.070, 0.071, 0.102, 0.030, 0.056, 0.148, 0.078, 0.089, 0.105, 0.039, 0.069, 0.120] rad 100 12-DOF Angle + 2-DOF Grasping 0.25 s 4 3×3×240×3203$\times$ 3$\times$ 240$\times$ 3203 × 3 × 240 × 320
Move the Stack 0.05 [0.049, 0.051, 0.057, 0.045, 0.038, 0.069] rad 100 6-DOF Angle + 1-DOF Grasping $\frac{1}{6}$ s 4 2×3×240×3202$\times$ 3$\times$ 240$\times$ 3202 × 3 × 240 × 320
TABLE IX: Summary of hyperparameters for trajectory optimization and dataset. $\alpha$ is the smoothing weight, $\frac{1}{\Delta t}$ represents the frequency of policy inference, and “Demos” indicates the number of demonstration episodes. “High Dim. Obs” refers to the dimensionality of high-dimensional observations. For instance, 2×3×84×842$\times$ 3$\times$ 84$\times$ 842 × 3 × 84 × 84 corresponds to RGB observations from two cameras, each with a shape of 3×84×843$\times$ 84$\times$ 843 × 84 × 84, while 3×5123$\times$ 5123 × 512 represents point cloud data with a shape of 3×5123$\times$ 5123 × 512. “Bound” refers to $d_{\text{max}}\Delta t$ in equation (7). “Unnormalized Bound” refers to constraint bounds unnormalized to dataset scale.
@@ VI-D Evaluating DiffOG under Multiple Motion Constraints
To further demonstrate the flexibility of DiffOG in handling diverse motion-related constraints, we conducted an experiment on the square task involving multiple motion constraints, as shown in Table X. In this new experiment, we additionally incorporated an explicit velocity constraint (with a normalized bound of 0.5) into the optimization problem (6), thereby enforcing both velocity and acceleration limits during training and inference. As shown in Table X, while the addition of a velocity constraint makes the feasible solution space strictly smaller and increases the difficulty of the optimization problem, DiffOG still achieves higher success rates than baseline methods that are only subject to the acceleration constraint. This highlights DiffOG’s flexibility in integrating multiple motion constraints and its effectiveness in producing trajectories that are both well-constrained and task-successful.
@@ VI-E Hyperparameters and Dataset
The summary of hyperparameters for trajectory optimization and dataset as shown in Table IX. In the experiments of this paper, we consistently adopt symmetric constraints which is $d_{\text{max}}=-d_{\text{min}}$. Since the actions require normalization to [-1,1] during training [5], all constraints are applied on the normalized action scale for DiffOG. For simplicity, when applying constraints on the normalized action scale, the same value is used for each action dimension. Therefore, the “Bound” column contains only a single value for each task. Since the distribution of action values differs across task datasets, the bounds for each action dimension may vary after unnormalization. For Robomimic tasks [13], the action values in the dataset are already normalized to the range of [1,1][-1,1][ - 1 , 1 ]. Therefore, the bound and the unnormalized bound are identical.
@@
The bound of 0.1 for Robomimic tasks represents normalized acceleration. To convert the it back to the raw form, use a scalar of 20 $\text{m}/\text{s}^{2}$ for linear acceleration and 200 $\text{rad}/\text{s}^{2}$ for second derivative of axis-angle [13]. For example, the maximum linear acceleration is $0.1\times 20~\text{m}/\text{s}^{2}=2~\text{m}/\text{s}^{2}$.
Since constraint clipping and penalty-based trajectory optimization do not require training, constraints can be directly applied to unnormalized actions. These constraints are equivalent to those imposed by DiffOG but are applied directly as unnormalized constraints. The parameters for penalty-based trajectory optimization are listed in Table XI. As mentioned earlier, the actions provided by the Robomimic dataset are already normalized to the range of [1,1][-1,1][ - 1 , 1 ]. The constraint bound we impose in this case is 0.10.10.1. Based on this, the violation tolerance for penalty-based optimization is set to a relatively small value of 0.0040.0040.004. For other tasks, the constraints are applied to raw action values, and the violation tolerances are determined based on the scale of unnormalized bounds provided in Table IX.

### v5

@@ IV-C Ablation Study on Static $\mathbf{Q}$, Matrix Learning, and Transformer
From equation (9), we can observe that if we set $\mathbf{Q}=-\mathbf{I}$, the term $\bigl\|\mathbf{R}\mathbf{y}_{t}-\mathbf{g}_{t}\bigr\|^{2}$ in optimization problem (8) simplifies to $\bigl\|\mathbf{y}_{t}-\mathbf{a}_{t}\bigr\|^{2}$, because $\mathbf{Q}=\mathbf{R}^{\mathrm{T}}\mathbf{R}$ and $\mathbf{g}_{t}=-(\mathbf{R}^{\mathrm{T}})^{-1}\mathbf{a}_{t}$. Under this setting, (8) becomes a trajectory optimization problem with both a smoothing cost and a distance cost. On the other hand, instead of using a transformer as in DiffOG, we can also directly learn the matrix $\mathbf{Q}$. In this section, we conduct experiments on the following three different cases:
@@
Additionally, during refine training, due to the presence of a multi-modal action distribution in demonstration dataset [chi2023diffusionpolicy, florence2022implicit], the actions generated by the pre-trained policy may differ from the targets in the dataset. This conflict between the targets and inputs does not exist in dataset training paradigm of DiffOG and matrix learning. From the results of matrix-learning refine and DiffOG refine, we can observe that trajectory optimization based on matrix learning struggles to handle such conflict. However, DiffOG can handle this conflict with transformer encoder.
@@
TABLE VI: We report the batched average inference time across three simulated tasks: lift, can, and square. Each measurement is based on a batch size of 25 and averaged over 50 inference runs. All methods are evaluated using 100 DDPM steps [ho2020denoising].
@@
Lastly, to demonstrate what the learned transformer and $\mathbf{Q}$ captures, we visualize its structure on the tool hang task. For each forward step, we compute a diagonality score defined as $\sum_{i}Q_{ii}^{2}/(\sum_{i,j}Q_{ij}^{2}+\epsilon)$, where $Q_{ij}$ denotes the $(i,j)$-th entry of the matrix $\mathbf{Q}$, and $\epsilon$ is a small constant (e.g., $1\times 10^{-4}$ ) added for numerical stability. A larger diagonality score indicates that $\mathbf{Q}$ is more diagonal, while a smaller value suggests stronger off-diagonal coupling. Fig. 10 shows that $\mathbf{Q}$ becomes more diagonal during fine-manipulation phases (B: grasping, D: insertion, F: grasping, H: hanging), and less diagonal during rapid transition phases. Intuitively, when the policy already outputs actions of low magnitude, a diagonal $\mathbf{Q}$ preserves these fine-grained action characteristics. In contrast, during fast movement phases, off diagonal elements allow the optimizer to capture correlations across different action dimensions and time steps, enabling smoother global trajectories. This example illustrates how DiffOG leverages data-driven structure in $\mathbf{Q}$ to retain task relevant policy behavior when appropriate and to inject coupling when necessary, thereby enhancing generalizability.
@@ IV-D Comparison with a Constrained Visuomotor Policy
We conducted comparison experiments between DiffOG and constrained visuomotor policy Leto [xu2024leto] on 7 tasks. The results are shown in Table VII. The results in the table compare DiffOG and Leto under the same constraint bounds.
@@
TABLE VII: Success rate of benchmarking DiffOG with Leto [xu2024leto].
@@ V Limitations and Future Work
Moreover, our method has demonstrated strong empirical performance under widely adopted action representations in imitation learning, such as joint angles [zhao2023learning, wu2024gello] and 6D end-effector velocities [mandlekar2022matters, wang2024poco, zhu2023viola]. Beyond these settings, we identify several promising directions for future work: 1) Certain action spaces that are less common in current imitation learning benchmarks, such as joint torques, also conform to the formulations of DiffOG. While not yet widely used, joint torques hold strong potential for force-based imitation learning, as they provide direct control over physical interactions. Extending our method to torque-based action spaces could enable more contact-rich manipulation. 2) Future extensions of the framework may address more complex representations, such as absolute $SO(3)$ rotations. In such cases, the assumption of simple discrete-time derivatives may no longer hold, requiring more expressive formulations of the underlying optimization problem.
@@ VI Conclusion
We introduced DiffOG, a learning-based trajectory optimization framework that combines the strengths of neural networks with an interpretable differentiable optimization layer. By integrating trajectory smoothness, adherence to hard constraints, and fidelity to demonstrations into the training objectives, DiffOG enables visuomotor policies to produce smoother and safer action trajectories while maintaining performance. Through extensive evaluations on 11 simulated tasks and 2 real-world tasks, we demonstrated that DiffOG consistently outperforms existing trajectory processing baselines, such as greedy constraint clipping and penalty-based trajectory optimization, as well as constrained visuomotor policy Leto [xu2024leto]. Our results demonstrate the generalization capability of DiffOG across diverse tasks.
@@ VI-B Action Space Assumption
1. The action space consists of joint angles, where the derivative corresponds to joint velocities. This type of action space is commonly used in systems that collect data through leader-follower arms, such as ALOHA [zhao2023learning] and GELLO [wu2024gello].
2. The action space consists of the robot end-effector linear position in xyzxyz, where the derivative corresponds to linear velocity. This action space is used in the push-T task in [chi2023diffusionpolicy] and the Meta-World benchmark [yu2020meta].
Another widely adopted action space is $v_{x},\,v_{y},\,v_{z},\,\dot{r}_{x},\,\dot{r}_{y},\,\dot{r}_{z}]^{\mathrm{T}}\in\mathbb{R}^{6},$ where $\mathbf{v}_{xyz}=[v_{x},v_{y},v_{z}]^{\mathrm{T}}$ is translational velocity, and $\dot{\mathbf{r}}=[\dot{r}_{x},\dot{r}_{y},\dot{r}_{z}]^{\mathrm{T}}$ is the derivative of the axis-angle rotation vector [mandlekar2022matters, wang2024poco, zhu2023viola, xu2024leto]. Its time derivative, $\dot{v}_{x},\,\dot{v}_{y},\,\dot{v}_{z},\,\ddot{r}_{x},\,\ddot{r}_{y},\,\ddot{r}_{z}]^{\!\mathrm{T}},$ includes linear accelerations and axis-angle second derivatives, which differ from the true angular acceleration $\dot{\bm{\omega}}$.
Although $\ddot{\mathbf{r}}$ does not directly represent $\dot{\bm{\omega}}$, constraining $\ddot{\mathbf{r}}$ effectively regulates $\dot{\bm{\omega}}$. The angular acceleration is related by $\dot{\bm{\omega}}=\mathbf{J}(\mathbf{r})\ddot{\mathbf{r}}+\dot{\mathbf{J}}(\mathbf{r})\dot{\mathbf{r}}$, where $\mathbf{J}(\mathbf{r})$ is the Jacobian matrix mapping axis-angle velocity to angular velocity. In practice, $\mathbf{J}(\mathbf{r})$ typically remains bounded, as evidenced by empirical successes of such action space in imitation learning frameworks [mandlekar2022matters, wang2024poco, zhu2023viola], which achieve stable rotations even without explicit constraints. Different from these prior work, DiffOG explicitly and effectively constrains rotational motion, resulting in smoother and more controlled behaviors. Extensive experimental results demonstrate the effectiveness of DiffOG.
@@ VI-C Task Descriptions and Experimental Setup
For the end-effector 6-DOF velocity action space, we include five tasks: lift, can, square, tool hang, and transport, all sourced from Robomimic benchmark [mandlekar2022matters]. For the end-effector linear position action space, we include six tasks: push-T from [chi2023diffusionpolicy], along with pick place wall, shelf place, disassemble, stick push, and stick pull from Meta-World [yu2020meta]. Notably, we selected all the “very hard” tasks identified in Meta-World as marked in [ze20243d]. Lastly, for the joint angle action space, we include two real-world tasks, arrange desk and move the stack, with data collection and policy experiments conducted using ALOHA [zhao2023learning].
@@
Lift 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 $2\times 3\times 84\times 84$
Can 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 $2\times 3\times 84\times 84$
Square 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 $2\times 3\times 84\times 84$
Tool Hang 0.1 0.1 200 6-DOF Vel. + 1-DOF Grasping 0.05 s 4 $2\times 3\times 240\times 240$
Transport 0.1 0.1 200 12-DOF Vel. + 2-DOF Grasping 0.05 s 4 $4\times 3\times 84\times 84$
Push-T 0.05 [12.5,12.2] Pixels 90 2-DOF Position 0.1 s 1 $1\times 3\times 96\times 96$
Pick Place Wall 0.6 [0.53, 0.92, 3.75] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Shelf Place 0.6 [2.32, 1.47, 4.13] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Disassemble 0.6 [0.65, 1.10, 0.61] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Stick Push 0.6 [1.03, 0.15, 0.68] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Stick Pull 0.6 [4.46, 1.44, 2.49] mm 30 3-DOF Pos. + 1-DOF Grasping 0.0125 s 4 $3\times 512$
Arrange Desk 0.1 [0.070, 0.071, 0.102, 0.030, 0.056, 0.148, 0.078, 0.089, 0.105, 0.039, 0.069, 0.120] rad 100 12-DOF Angle + 2-DOF Grasping 0.25 s 4 $3\times 3\times 240\times 320$
Move the Stack 0.05 [0.049, 0.051, 0.057, 0.045, 0.038, 0.069] rad 100 6-DOF Angle + 1-DOF Grasping $\frac{1}{6}$ s 4 $2\times 3\times 240\times 320$
TABLE IX: Summary of hyperparameters for trajectory optimization and dataset. $\alpha$ is the smoothing weight, $\frac{1}{\Delta t}$ represents the frequency of policy inference, and “Demos” indicates the number of demonstration episodes. “High Dim. Obs” refers to the dimensionality of high-dimensional observations. For instance, $2\times 3\times 84\times 84$ corresponds to RGB observations from two cameras, each with a shape of $3\times 84\times 84$, while $3\times 512$ represents point cloud data with a shape of $3\times 512$. “Bound” refers to $d_{\text{max}}\Delta t$ in equation (7). “Unnormalized Bound” refers to constraint bounds unnormalized to dataset scale.
@@ VI-D Evaluating DiffOG under Multiple Motion Constraints
To further demonstrate the flexibility of DiffOG in handling diverse motion-related constraints, we conducted an experiment on the square task involving multiple motion constraints, as shown in Table X. In this new experiment, we additionally incorporated an explicit velocity constraint (with a normalized bound of 0.5) into optimization problem (6), thereby enforcing both velocity and acceleration limits during training and inference. As shown in Table X, while the addition of a velocity constraint makes the feasible solution space strictly smaller and increases the difficulty of the optimization problem, DiffOG still achieves higher success rates than baseline methods that are only subject to the acceleration constraint. This highlights DiffOG’s flexibility in integrating multiple motion constraints and its effectiveness in producing trajectories that are both well-constrained and task-successful.
@@ VI-E Hyperparameters and Dataset
The summary of hyperparameters for trajectory optimization and dataset as shown in Table IX. In the experiments of this paper, we consistently adopt symmetric constraints which is $d_{\text{max}}=-d_{\text{min}}$. Since the actions require normalization to [-1,1] during training [chi2023diffusionpolicy], all constraints are applied on the normalized action scale for DiffOG. For simplicity, when applying constraints on the normalized action scale, the same value is used for each action dimension. Therefore, the “Bound” column contains only a single value for each task. Since the distribution of action values differs across task datasets, the bounds for each action dimension may vary after unnormalization. For Robomimic tasks [mandlekar2022matters], the action values in the dataset are already normalized to the range of [1,1][-1,1]. Therefore, the bound and the unnormalized bound are identical.
@@
The bound of 0.1 for Robomimic tasks represents normalized acceleration. To convert the it back to the raw form, use a scalar of 20 $\text{m}/\text{s}^{2}$ for linear acceleration and 200 $\text{rad}/\text{s}^{2}$ for second derivative of axis-angle [mandlekar2022matters]. For example, the maximum linear acceleration is $0.1\times 20~\text{m}/\text{s}^{2}=2~\text{m}/\text{s}^{2}$.
Since constraint clipping and penalty-based trajectory optimization do not require training, constraints can be directly applied to unnormalized actions. These constraints are equivalent to those imposed by DiffOG but are applied directly as unnormalized constraints. The parameters for penalty-based trajectory optimization are listed in Table XI. As mentioned earlier, the actions provided by the Robomimic dataset are already normalized to the range of [1,1][-1,1]. The constraint bound we impose in this case is $0.1$. Based on this, the violation tolerance for penalty-based optimization is set to a relatively small value of $0.004$. For other tasks, the constraints are applied to raw action values, and the violation tolerances are determined based on the scale of unnormalized bounds provided in Table IX.